OpenStudy (anonymous):

When the number of marbles Jill has is subtracted from 20, the result is the same as adding 16 to the number of marbles. How many marbles does Jill have? So is it 20 - 16 = 4 + 16

OpenStudy (goformit100):

@Miracrown

OpenStudy (anonymous):

Can you help me with this one @micracrown?

OpenStudy (anonymous):

@Miracrown

OpenStudy (alekos):

20-x=x+16

OpenStudy (alekos):

Solve for x

OpenStudy (anonymous):

would x = 4

OpenStudy (anonymous):

20 + 16 = 36 as +16 +16 = 32 +4

OpenStudy (alekos):

X=2